OpenStudy (anonymous):

differentiate using quotient rule: s^1/3 - 1 / s^2/3 - 1

OpenStudy (anonymous):

I presume you mean \[(s^(1\div3)-1)\div(s^(2\div3)-1)\] sorry, this looks a bit messy. In words "(cube root of s, minus1) on top line, (cube root of s all squared, minus 1) on the bottom line" If so, bottom line is a difference of two squares, so simplify forst before differentiating

OpenStudy (anonymous):

how is the bottom line the difference of two squares?

OpenStudy (anonymous):

do you mean (s^1/3)^2 - (1)^2

OpenStudy (anonymous):

Difference of two squares is like a^2 - b^2 In this case a = s^(1/3) and b = 1, so bottom line is (S^(1/3) -1)(s^(1/3) + 1) then cancel top and bottom factors to get a much easier expression to differentiate.

OpenStudy (anonymous):

oh yeah.. that made it a lot easier. thanks! :)